two cars each moving at constant speed on a straight road pass a certain point on the road at the same time if car A is moving at 23 m/s and car B is moving at 14 m/s how far apart will they be after five minutes if they contains at the same speed?


1
Expert's answer
2022-03-14T15:14:13-0400

Distance = speed * time

5 minutes = 5 * 60 = 300 seconds

for car 1;

distance = 23 * 300 = 6900m

for car 2;

distance = 14 * 300 = 4200m

distance between them = 6900 - 4200

= 2700m